# Cross-Domain Few-Shot Object Detection via Enhanced Open-Set Object Detector

### v2

@@ Abstract
This paper studies the challenging cross-domain few-shot object detection (CD-FSOD), aiming to develop an accurate object detector for novel domains with minimal labeled examples. While transformer-based open-set detectors, such as DE-ViT, show promise in traditional few-shot object detection, their generalization to CD-FSOD remains unclear: 1) can such open-set detection methods easily generalize to CD-FSOD? 2) If not, how can models be enhanced when facing huge domain gaps? To answer the first question, we employ measures including style, inter-class variance (ICV), and indefinable boundaries (IB) to understand the domain gap. Based on these measures, we establish a new benchmark named CD-FSOD to evaluate object detection methods, revealing that most of the current approaches fail to generalize across domains. Technically, we observe that the performance decline is associated with our proposed measures: style, ICV, and IB. Consequently, we propose several novel modules to address these issues. First, the learnable instance features align initial fixed instances with target categories, enhancing feature distinctiveness. Second, the instance reweighting module assigns higher importance to high-quality instances with slight IB. Third, the domain prompter encourages features resilient to different styles by synthesizing imaginary domains without altering semantic contents. These techniques collectively contribute to the development of the Cross-Domain Vision Transformer for CD-FSOD (CD-ViTO), significantly improving upon the base DE-ViT. Experimental results validate the efficacy of our model. All datasets, codes, and models will be released to the community.
@@ 1 Introduction
Figure 1: (a) Our motivation: The DE-ViT open-set detector excels in FSOD but struggles in CD-FSOD, inspiring our creation of CD-ViTO. (b) Technical motivation: FSOD models face challenges when dealing with cross-domain targets, such as small inter-class variance, undefined boundaries, and varying appearances.
To address question ①, we establish a new benchmark for CD-FSOD by rearranging existing object detection datasets. Our benchmark utilizes COCO [29] as source training set and incorporates six additional datasets, namely ArTaxOr [2], Clipart1K [18], DIOR [26], DeepFish [36], NEU-DET [37], and UODD [20], as novel target datasets. We also introduce metrics such as style, Inter-Class Variance (ICV), and Indefinable Boundaries (IB) to measure domain differences. Style represents different visual appearances, ICV measures dissimilarity between semantic labels, and IB reflects confusion between targets and backgrounds. Our target datasets exhibit significant variations in style, ICV, and IB, making our benchmark a comprehensive testbed for evaluating CD-FSOD methods. We evaluate various object detectors on this benchmark. As shown in Fig. 1(a), we observe obvious performance degradation even for the State-Of-The-Art (SOTA) FSOD detector, DE-ViT. This indicates the challenge of tackling CD-FSOD even with well-trained open-set detectors.
To tackle both the performance degradation issue, i.e., question ②, we aim to upgrade the vanilla DE-ViT to improve the open-set detector’s performance. Our technical motivation, illustrated in Fig. 1(b), highlights that FSOD model degradation correlates with domain metrics we mentioned earlier: style, ICV, and IB. Specifically, a model trained on instances with coarse semantic concepts (large ICV) and clear foreground-background boundaries (slight IB), like "cup" and "cat", can easily transfer to similar in-domain concepts like "bird" and "dog". However, cross-domain cases, such as industrial defect types like "inclusion" and "patches", pose challenges with smaller ICV, more significant IB, and different styles. This insight inspires us to design a method that addresses issues of small ICV, significant IB, and changing styles in cross-domain scenarios.
@@
Furthermore, we also explore the finetuning technique. Although finetuning has been studied in cross-domain tasks [12, 6, 28], we provide a detailed analysis of its effectiveness for our benchmark, particularly in the context of open-set detectors. By integrating these novel modules into DE-ViT, we introduce a new vision transformer based open-set detector called CD-ViTO for CD-FSOD. These modifications substantially boost DE-ViT’s performance across all target datasets, as shown in Fig. 1(a) (highlighted by the orange stars).
We summarize our contributions. 1) We propose a comprehensive benchmark for CD-FSOD that spans diverse target domains. Additionally, we introduce several metrics related to domain gap analysis, including style, inter-class variance, and indefinable boundaries. 2) Through extensive evaluations, we uncover the challenges posed by the domain gap issue in existing open-set object detectors and other object detection methods for CD-FSL. 3) Leveraging the existing DE-ViT, we present CD-ViTO, a novel method that incorporates finetuning, learnable instance features, an instance reweighting module, and a domain prompter. Notably, these new modules introduce minimal parameters and require only a few target support images for training, ensuring high efficiency. 4) Our results demonstrate that CD-ViTO significantly improves upon the base DE-ViT, achieving new state-of-the-art performance in CD-FSOD.
@@ 2 Related Works
Few-Shot Object Detection. FSOD [23] identifies objects with few labeled data. Existing FSOD methods can be roughly divided into meta-learning based approaches [46, 13, 15, 21, 14] and transfer learning-based methods [42, 38, 4, 33, 31, 22, 11]. Meta-learning trains model across tasks for quick adaptation. Meta-RCNN [46] is a typical work of this type. Transfer learning finetunes pretrained model on limited data for the few-shot task. Flagship works include TFA [42], FSCE [38], and DeFRCN [33]. Though methods may propose specific techniques e.g., contrastive encoding in FSCE, and decoupled framework in DeFRCN, generally they still fall into either the meta-learning based or finetuning based types. A recently proposed method, DE-ViT [49], unifying the problem of FSOD and open-vocabulary object detection [47] and constructing an open-set detector for both tasks, shows great advantage. However, almost all of the methods mentioned above ideally assume that the training and testing sets belong to the same domain.
Exceptions include Distill-cdfsod [44] and MoFSOD [24]. Distill-cdfsod [44] introduces several datasets and a distillation-based method, but both the benchmark and method have room for improvement. MoFSOD [24] also explores CD-FSOD and studies various model architectures, finetuning, and pre-trained datasets. However, apart from the datasets, we for the first time explore the problem from an open-set perspective with a novel algorithm introduced. Note that several methods, such as AcroFOD [9] and AsyFOD [8], also explore FSOD across domains. However, we are essentially addressing different problems. They share the same class set between the source and target domains, while we have to handle entirely novel target categories.
Cross-Domain Few-Shot Learning. CD-FSL [45] aims to improve few-shot learning (FSL) across diverse domains. While there are several existing works [39, 12, 40, 5, 54, 28, 17, 6] addressing classification tasks, there is a noticeable gap in tackling domain gap challenges across various few-shot vision tasks, including object detection [46, 13], semantic segmentation [41, 43], and action recognition [48, 7]. This paper specifically explores the area of CD-FSOD, seeking to develop a robust object detector. Despite recent advancements [9, 44, 24] in CD-FSOD, certain crucial aspects, such as domain gap metrics, comprehensive baseline evaluations, and technical improvements, remain less explored. Thus, we meticulously organize the benchmark, conduct an in-depth study on various existing methods, and introduce CD-ViTO as a novel method. Our goal is to propel the field of CD-FSOD forward, shedding light on unexplored aspects.
@@ 3.1 Task Configurations
Training Strategy. We adopt a "pretrain, finetune, and testing" pipeline, as used in many previous CD-FSL methods [12, 17]. This pipeline involves training the evaluated methods on the source dataset $\mathcal{D}_{S}$, followed by finetuning the trainable parameters using the few labeled support images $S$ from specific target dataset $\mathcal{D}_{T}$, and finally testing the model on the novel query set $Q$. Typical objective functions for training and finetuning include the box regression loss and the classification loss.
@@ 4.1 Overview of CD-ViTO
Preliminary. DE-ViT  [49] builds an open-set detector by using the visual features from large pretrained model and disentangling the localization and classification tasks. The pipeline for the base DE-ViT is represented by blue arrows in Fig. 3(a). It mainly contains a pretrained DINOv2 ViT [32], a region proposal network ($M_{RPN}$), an ROI align module ($M_{ROI}$), a detection head ($M_{DET}$), and a one-vs-rest classification head ($M_{CLS}$). Particularly, given a query image $q$ and a set of support images $S$, DE-ViT first uses the DINOv2 to extract the instance features $F_{ins}=\{{F_{ins}^{ob},F_{ins}^{bg}}\}$, where the $F_{ins}^{ob}$ represents features for foreground objects from $S$ and the $F_{ins}^{bg}$ denotes features for background instances. Then, the class prototypes $F_{pro}=\{{F_{pro}^{ob},F_{ins}^{bg}}\}$ are obtained by averaging $F_{ins}^{ob}$ on object categories while keeping $F_{ins}^{bg}$ unchanged (as in blue dashed arrows). For the query image $q$, DE-ViT applies the DINOv2, $M_{RPN}$, and $M_{ROI}$ to generate the region proposals $R_{q}$, visual features $F_{q}$, and ROI features $F_{q_{roi}}$. After that, the $M_{DET}$ takes the $R_{q}$, $F_{q}$, and $F_{pro}$ as input for the localization task producing $\mathcal{L}_{loc}$. In the meantime, the $M_{CLS}$ performs the classification task based on $F_{q_{roi}}$ and $F_{pro}$ resulting in $\mathcal{L}_{cls}$. The network is optimized by $\mathcal{L}_{loc}$ and $\mathcal{L}_{cls}$.
@@
Despite DE-ViT’s purported ability to detect any novel objects without finetuning, our experiments reveal its failure to generalize to cross-domain target datasets (Sec. 5.2). This motivates us to improve the cross-domain ability of DE-ViT and form a new method. Note that the basic components of DE-ViT including the DINOv2, $M_{RPN}$, $M_{ROI}$, $M_{DET}$, $M_{CLS}$, $\mathcal{L}_{loc}$, and $\mathcal{L}_{cls}$, validated on FSOD tasks, are directly inherited in our framework. Our method, CD-ViTO, is outlined in Fig. 3(a) with new enhancements highlighted in orange. Recall that in Sec. 1, we point out that the domain gap of CD-FSOD is highly related to the ICV, IB, and style of the target datasets. Therefore, we are inspired to design our CD-ViTO by tackling these factors. As a result, the novel learnable instance features ($M_{LIF}$), the instance reweighting ($M_{LR}$), the domain prompter ($M_{DP}$) are proposed to tackle the small ICV, significant IB, and changing styles, respectively. In addition, though being explored in related cross-domain tasks [12, 6, 28], the finetuning (FT) is also explored in our benchmark.
@@
Specifically, given $S$ and $q$ as input, we perform the same steps as DE-ViT to obtain instance features $F_{ins}=\{F_{ins}^{ob},F_{ins}^{bg}\}$, region proposals $R_{q}$, visual features $F_{q}$, and ROI features $F_{q_{roi}}$. But different from directly averaging $F_{ins}^{ob}$ to obtain class prototypes $F_{pro}^{ob}$ for objects, we first utilize our $M_{LIF}$ to make the precalculated $F_{ins}$ learnable and adapt them by performing the disentangled detection tasks i.e., the localization task supervised by $\mathcal{L}_{loc}$ and the classification task supervised by $\mathcal{L}_{cls}$. Since the training data is labeled with class categories, such an optimization would align the instance features to their corresponding categories thus increasing the discriminability of features and addressing the small ICV issue. Then, considering the qualities especially the IB of object instances varies, we apply our $M_{LR}$ to reweight the weights of object instances features $F_{ins}^{ob}$. Generally, $M_{LR}$ is expected to assign higher values to the high-quality instances with slight IB. In this way, the high-quality instances could contribute more to its class prototype. The $M_{LR}$ is inserted after the $M_{LIF}$ and is optimized use the same $\mathcal{L}_{loc}$, $\mathcal{L}_{cls}$. In addition, to make the model more robust to the changing styles, our proposed $M_{DP}$ is applied to synthesize several virtual "domains" $F_{domain}$. The goal of the $M_{DP}$ is to introduce such domain perturbations and force our features to be consistent with different domains. To ensure the diversity of the introduced domains and constrain the domains would not affect the semantics of features, an extra loss $\mathcal{L}_{dp}$ is introduced. The $M_{DP}$ is build upon the object prototypes $F_{pro}^{ob}$ and the $\mathcal{L}_{dp}$ works with the $\mathcal{L}_{loc}$ and $\mathcal{L}_{cls}$ together.
@@ 4.2 Key Components of Our CD-ViTO
In this subsection, we give a detailed introduction to our novel components $M_{LIF}$, $M_{IR}$, and $M_{DP}$.
@@
Instance Reweighting Module $M_{IR}$. Our $M_{IR}$ is proposed to reweight the instances enabling high-quality object instances e.g., objects with slight IB contribute more to the class prototype. The architecture of $M_{IR}$ is shown in Fig. 3(b). Given the learnable instance features $F_{ins}^{lea}$ from $M_{LIF}$, we first obtain the object instance features from $F_{ins}^{lea}$ and denote it as $F_{ins}^{0}\in\mathcal{R}^{N\times K\times D}$. Our $M_{IR}$ has two paths connected in a residual way. More specifically, the lower path feeds the $F_{ins}^{0}$ into an MLP module to get the weighted score which is further used to perform the weighted sum on the initial $F_{ins}^{0}$ resulting in the deformed prototypes $F_{ins}^{att}\in\mathcal{R}^{N\times D}$. The upper path performs the average operation on $F_{ins}^{0}$ getting $F_{ins}^{avg}\in\mathcal{R}^{N\times D}$. The final object prototypes $F_{pro}^{ob}$ are represented as $\alpha fc(F_{ins}^{att})+(1-\alpha)F_{ins}^{avg}$, where the $fc(\cdot)$ denotes a fully connected layer and $\alpha$ is a hyperparameter.
Domain Prompter $M_{DP}$. To make our object prototype features $F_{pro}^{ob}$ robust to various domains, we have to solve two goals: 1) synthesizing several "domains" with diversities; 2) constraining that perturbing features with the synthesized domains will not cause the semantic drift. To achieve these, our $M_{DP}$ first introduces several learnable vectors $F_{domain}\in\mathcal{R}^{N_{dom}\times D}$ as virtual domains and then proposes several loss functions as supervision. The concept of "prompter" is also related but differs from the prompt tuning [19, 25]. Notably, both approaches employ trainable parameters as prompters, but general prompt tuning facilitates the transfer of pretrained models through efficient tuning, whereas we introduce prompters as perturbations. The illustration of our $M_{DP}$ is demonstrated in Fig. 3(b). $N_{dom}$ is a hyperparameter.
Concretely, given $F_{domain}$ and $F_{pro}^{ob}$, we first propose the domain diversity loss $\mathcal{L}_{domain}$ to force the different domains e.g., $f^{d_{i}}$ and $f^{d_{j}}$ to be far from each other. Secondly, for a prototype $f_{p_{i}}$ in $F_{pro}^{ob}$, we randomly perturbs it with two various domains $f^{d_{k}}$ and $f^{d_{m}}$ sampled from $F_{domain}$ forming two perturbed prototypes $f_{p_{i}}^{d_{k}},f_{p_{i}}^{d_{m}}$. The perturbation is done by adding the features e.g., $f_{p_{i}}^{d_{k}}=f_{p_{i}}+f^{d_{k}}$. The domain $f^{d_{m}}$ is also added to other object prototypes forming $f_{p_{j}}^{d_{m}}$, where $j=1,\ldots,N$. We constrain that $f_{p_{i}}^{d_{k}}$ and $f_{p_{i}}^{d_{m}}$ should be close to each other while being distant from perturbed prototypes generated from different class prototypes. This constitutes our prototype consistency loss $\mathcal{L}_{proto}$. Both the $\mathcal{L}_{domain}$ and $\mathcal{L}_{proto}$ are implemented by the InfoNCE loss as follows:

### v3

@@ Abstract
This paper studies the challenging cross-domain few-shot object detection (CD-FSOD), aiming to develop an accurate object detector for novel domains with minimal labeled examples. While transformer-based open-set detectors, such as DE-ViT, show promise in traditional few-shot object detection, their generalization to CD-FSOD remains unclear: 1) can such open-set detection methods easily generalize to CD-FSOD? 2) If not, how can models be enhanced when facing huge domain gaps? To answer the first question, we employ measures including style, inter-class variance (ICV), and indefinable boundaries (IB) to understand the domain gap. Based on these measures, we establish a new benchmark named CD-FSOD to evaluate object detection methods, revealing that most of the current approaches fail to generalize across domains. Technically, we observe that the performance decline is associated with our proposed measures: style, ICV, and IB. Consequently, we propose several novel modules to address these issues. First, the learnable instance features align initial fixed instances with target categories, enhancing feature distinctiveness. Second, the instance reweighting module assigns higher importance to high-quality instances with slight IB. Third, the domain prompter encourages features resilient to different styles by synthesizing imaginary domains without altering semantic contents. These techniques collectively contribute to the development of the Cross-Domain Vision Transformer for CD-FSOD (CD-ViTO), significantly improving upon the base DE-ViT. Experimental results validate the efficacy of our model. Datasets and codes are available at http://yuqianfu.com/CDFSOD-benchmark.
@@ 1 Introduction
Figure 1: (a) Our motivation: The DE-ViT open-set detector excels in FSOD but struggles in CD-FSOD, inspiring our creation of CD-ViTO. (b) Technical motivation: FSOD models face challenges when dealing with cross-domain targets, such as small inter-class variance (ICV), indefinable boundaries (IB), and varying appearances (styles).
To address question ①, we establish a new benchmark for CD-FSOD by rearranging existing object detection datasets. Our benchmark utilizes COCO [28] as source training set and incorporates six additional datasets, namely ArTaxOr [2], Clipart1K [18], DIOR [26], DeepFish [36], NEU-DET [38], and UODD [20], as novel target datasets. We also introduce metrics such as style, Inter-Class Variance (ICV), and Indefinable Boundaries (IB) to measure domain differences. Style represents different visual appearances, ICV measures dissimilarity between semantic labels, and IB reflects confusion between targets and backgrounds. Our target datasets exhibit significant variations in style, ICV, and IB, making our benchmark a comprehensive testbed for evaluating CD-FSOD methods. We evaluate various object detectors on this benchmark. As shown in Fig. 1(a), we observe obvious performance degradation even for the State-Of-The-Art (SOTA) FSOD detector, DE-ViT. This indicates the challenge of tackling CD-FSOD even with well-designed open-set detectors.
To tackle the performance degradation issue, i.e., question ②, we aim to upgrade the vanilla DE-ViT to improve the open-set detector’s performance. Our technical motivation, illustrated in Fig. 1(b), highlights that FSOD model degradation correlates with domain metrics we mentioned earlier: style, ICV, and IB. Specifically, a model trained on instances with coarse semantic concepts (large ICV) and clear foreground-background boundaries (slight IB), like "cup" and "cat", can easily transfer to similar in-domain concepts like "bird" and "dog". However, cross-domain cases, such as industrial defect types like "inclusion" and "patches", pose challenges with smaller ICV, more significant IB, and different styles. This insight inspires us to design a method that addresses issues of small ICV, significant IB, and changing styles in cross-domain scenarios.
@@
Furthermore, we also explore the finetuning technique. Although finetuning has been studied in cross-domain tasks [12, 6], we provide a detailed analysis of its effectiveness on our benchmark. By integrating these novel modules into DE-ViT, we introduce a new vision transformer based open-set detector called CD-ViTO for CD-FSOD. CD-ViTO substantially boosts DE-ViT’s performance across all target datasets, as shown in Fig. 1(a) (highlighted by the orange stars).
We summarize our contributions. 1) We propose a comprehensive benchmark for CD-FSOD that spans diverse target domains. Additionally, we introduce several metrics related to domain gap analysis, including style, inter-class variance, and indefinable boundaries. 2) Through extensive evaluations, we uncover the challenges posed by the domain gap issue in existing open-set object detectors and other object detection methods for CD-FSL. 3) Leveraging the existing DE-ViT, we present CD-ViTO, a novel method that incorporates finetuning, learnable instance features, an instance reweighting module, and a domain prompter. Notably, these new modules introduce minimal parameters and require only a few target support instances for training, ensuring high efficiency. 4) Our results demonstrate that CD-ViTO significantly improves upon the base DE-ViT, achieving new state-of-the-art performance in CD-FSOD.
@@ 2 Related Works
Few-Shot Object Detection. FSOD [23] identifies objects with few labeled data. Existing FSOD methods can be roughly divided into meta-learning based [48, 13, 15, 21, 14] and transfer learning based [45, 39, 4, 33, 31, 22, 11]. Meta-learning trains model across tasks for quick adaptation. Meta-RCNN [48] is a typical work of this type. Transfer learning finetunes pretrained model on limited data for few-shot tasks. Flagship works include TFA [45], FSCE [39], and DeFRCN [33]. Notably, a recently proposed method, DE-ViT [52], unifying the problem of FSOD and open-vocabulary object detection [49] and constructing an open-set detector for both tasks, shows great advantage. However, almost all of the methods mentioned above ideally assume that the training and testing sets belong to the same domain. Exceptions include Distill-cdfsod [47] and MoFSOD [24]. Distill-cdfsod [47] introduces several datasets and a distillation-based method, but both the benchmark and method have room for improvement. MoFSOD [24] also explores CD-FSOD and studies various model architectures, finetuning, and pre-trained datasets. Compared with them, apart from the datasets, we for the first time propose benchmark metrics and study the open-set detectors for CD-FSOD. Note that AcroFOD [9] and AsyFOD [8] also explore FSOD across domains. However, they share the same class set between source and target domains, while we handle entirely novel target categories.
Cross-Domain Few-Shot Learning. CD-FSL aims to improve few-shot learning (FSL) [37, 42, 51, 29, 40] across diverse domains. While there are several existing works [41, 12, 43, 5, 57, 17, 6] addressing classification tasks, there is a noticeable gap in tackling domain gap challenges across various few-shot vision tasks, including object detection [48, 13], semantic segmentation [44, 46], and action recognition [50, 7]. This paper specifically explores the area of CD-FSOD, seeking to develop a robust object detector. Despite recent advancements [9, 47, 24] in CD-FSOD, certain crucial aspects, such as domain gap metrics, comprehensive baseline evaluations, and technical improvements, remain less explored. Thus, we meticulously organize the benchmark, conduct an in-depth study on various existing methods, and introduce CD-ViTO as a novel method. Our goal is to propel the field of CD-FSOD forward, shedding light on unexplored aspects.
@@ 3.1 Task Configurations
Training Strategy. We adopt a "pretrain, finetune, and testing" pipeline, as used in many previous CD-FSL methods [12, 17]. This pipeline involves training methods on the source dataset $\mathcal{D}_{S}$, followed by finetuning the trainable parameters using the few labeled support instances $S$ from target dataset $\mathcal{D}_{T}$, and finally testing on the novel query set $Q$. Typical objective functions for training and finetuning include the box regression loss and the classification loss.
@@ 4.1 Overview of CD-ViTO
Preliminary. DE-ViT  [52] builds an open-set detector by using the visual features from large pretrained model and disentangling the localization and classification tasks. The pipeline for the base DE-ViT is represented by blue arrows in Fig. 3(a). It mainly contains a pretrained DINOv2 ViT [32], a region proposal network ($M_{RPN}$), an ROI align module ($M_{ROI}$), a detection head ($M_{DET}$), and a one-vs-rest classification head ($M_{CLS}$). Particularly, given a query image $q$ and a set of support instances $S$, DE-ViT first uses the DINOv2 to extract the instance features $F_{ins}=\{{F_{ins}^{ob},F_{ins}^{bg}}\}$, where the $F_{ins}^{ob}$ represents features for foreground objects from $S$ and the $F_{ins}^{bg}$ denotes features for background instances. Then, the class prototypes $F_{pro}=\{{F_{pro}^{ob},F_{ins}^{bg}}\}$ are obtained by averaging $F_{ins}^{ob}$ on object categories while keeping $F_{ins}^{bg}$ unchanged (as in blue dashed arrows). For the query image $q$, DE-ViT applies the DINOv2, $M_{RPN}$, and $M_{ROI}$ to generate the region proposals $R_{q}$, visual features $F_{q}$, and ROI features $F_{q_{roi}}$. After that, the $M_{DET}$ takes the $R_{q}$, $F_{q}$, and $F_{pro}$ as input for the localization task producing $\mathcal{L}_{loc}$. In the meantime, the $M_{CLS}$ performs the classification task based on $F_{q_{roi}}$ and $F_{pro}$ resulting in $\mathcal{L}_{cls}$. The network is optimized by $\mathcal{L}_{loc}$ and $\mathcal{L}_{cls}$.
@@
Despite DE-ViT’s purported ability to detect any novel objects without finetuning, our experiments reveal its failure to generalize to cross-domain target datasets (Sec. 5.2). This motivates us to improve the cross-domain ability of DE-ViT and form a new method. Note that the basic components of DE-ViT including the DINOv2, $M_{RPN}$, $M_{ROI}$, $M_{DET}$, $M_{CLS}$, $\mathcal{L}_{loc}$, and $\mathcal{L}_{cls}$, validated on FSOD tasks, are directly inherited in our framework. Our method, CD-ViTO, is outlined in Fig. 3(a) with new enhancements highlighted in orange. Recall that in Sec. 1, we point out that the domain gap of CD-FSOD is highly related to the ICV, IB, and style of the target datasets. Therefore, we are inspired to design our CD-ViTO by tackling these factors. As a result, the novel learnable instance features ($M_{LIF}$), the instance reweighting ($M_{IR}$), the domain prompter ($M_{DP}$) are proposed to tackle the small ICV, significant IB, and changing styles, respectively. In addition, though being explored in related cross-domain tasks [12, 6], the finetuning (FT) is also explored in our benchmark.
@@
Specifically, given $S$ and $q$ as input, we perform the same steps as DE-ViT to obtain instance features $F_{ins}=\{F_{ins}^{ob},F_{ins}^{bg}\}$, region proposals $R_{q}$, visual features $F_{q}$, and ROI features $F_{q_{roi}}$. But different from directly averaging $F_{ins}^{ob}$ to obtain class prototypes $F_{pro}^{ob}$ for objects, we first utilize our $M_{LIF}$ to make the precalculated $F_{ins}$ learnable and adapt them as $F_{ins}^{lea}$ by performing the disentangled detection tasks i.e., the localization task supervised by $\mathcal{L}_{loc}$ and the classification task supervised by $\mathcal{L}_{cls}$. Since the training data is labeled with class categories, such an optimization would align the instance features to their corresponding categories thus increasing the discriminability of features and addressing the small ICV issue. Then, considering the qualities especially the IB of object instances varies, we apply our $M_{IR}$ to reweight the weights of objects in the $F_{ins}^{lea}$, this step forms the new prototypes $F_{pro}^{{}^{\prime}}$. Generally, $M_{IR}$ is expected to assign higher values to the high-quality instances with slight IB. In this way, the high-quality instances could contribute more to its class prototype. The $M_{IR}$ is inserted after the $M_{LIF}$ and is optimized use the same $\mathcal{L}_{loc}$, $\mathcal{L}_{cls}$. In addition, to make the model more robust to the changing styles, our proposed $M_{DP}$ is applied to synthesize several virtual "domains" $F_{domain}$. The goal of the $M_{DP}$ is to introduce such domain perturbations and force our features to be consistent with different domains. To ensure the diversity of the introduced domains and constrain the domains would not affect the semantics of features, an extra loss $\mathcal{L}_{dp}$ is introduced. The $M_{DP}$ is build upon the prototypes reweighed by $M_{IR}$ i.e., $F_{pro}^{{}^{\prime}}$, and the $\mathcal{L}_{dp}$ works with the $\mathcal{L}_{loc}$ and $\mathcal{L}_{cls}$ together.
@@ 4.2 Key Components of Our CD-ViTO
Instance Reweighting Module $M_{IR}$. We propose $M_{IR}$ to enable high-quality object instances e.g., objects with slight IB contribute more to the class prototype. The architecture of $M_{IR}$ is shown in Fig. 3(b). Given the learnable instance features $F_{ins}^{lea}$ from $M_{LIF}$, we first obtain the object instance features from $F_{ins}^{lea}$ and denote it as $F_{ins}^{0}\in\mathcal{R}^{N\times K\times D}$. Our $M_{IR}$ has two paths connected in a residual way. More specifically, the lower path feeds $F_{ins}^{0}$ into an MLP module to get the weighted score which is further used to perform the weighted sum on the initial $F_{ins}^{0}$ resulting in the deformed prototypes $F_{pro}^{att}\in\mathcal{R}^{N\times D}$. The upper path performs the average operation on $F_{ins}^{0}$ getting $F_{pro}^{avg}\in\mathcal{R}^{N\times D}$. The final object prototypes $F_{pro}^{ob^{\prime}}$ are represented as $\alpha fc(F_{pro}^{att})+(1-\alpha)F_{pro}^{avg}$, where $fc(\cdot)$ is a fully connected layer and $\alpha$ is a hyperparameter. With $F_{pro}^{ob^{\prime}}$ and background features, we form new prototype $F_{pro}^{{}^{\prime}}$.
Domain Prompter $M_{DP}$. To make our object prototype features $F_{pro}^{ob^{\prime}}$ robust to various domains, we have to solve two goals: 1) synthesize several "domains" with diversities; 2) prevent semantic drift when perturbing features with such "domains". To achieve these, our $M_{DP}$ first introduces several learnable vectors $F_{domain}\in\mathcal{R}^{N_{dom}\times D}$ as virtual domains and then proposes several loss functions as supervision. The concept of "prompter" is also related but differs from the prompt tuning [19, 25]. Notably, both approaches employ trainable parameters as prompters, but general prompt tuning facilitates the transfer of pretrained models through efficient tuning, whereas we introduce prompters as perturbations. The illustration of $M_{DP}$ is demonstrated in Fig. 3(b). $N_{dom}$ is a hyperparameter.
Concretely, given $F_{domain}$ and $F_{pro}^{ob^{\prime}}$, we first propose the domain diversity loss $\mathcal{L}_{domain}$ to force the different domains e.g., $f^{d_{i}}$ and $f^{d_{j}}$ to be far from each other. Secondly, for a prototype $f_{p_{i}}$ in $F_{pro}^{ob^{\prime}}$, we randomly perturbs it with two various domains $f^{d_{k}}$ and $f^{d_{m}}$ sampled from $F_{domain}$ forming two perturbed prototypes $f_{p_{i}}^{d_{k}},f_{p_{i}}^{d_{m}}$. The perturbation is done by adding the features e.g., $f_{p_{i}}^{d_{k}}=f_{p_{i}}+f^{d_{k}}$. The domain $f^{d_{m}}$ is also added to other object prototypes forming $f_{p_{j}}^{d_{m}}$, where $j=1,\ldots,N$. We constrain that $f_{p_{i}}^{d_{k}}$ and $f_{p_{i}}^{d_{m}}$ should be close to each other while being distant from perturbed prototypes generated from different class prototypes. This constitutes our prototype consistency loss $\mathcal{L}_{proto}$. Both the $\mathcal{L}_{domain}$ and $\mathcal{L}_{proto}$ are implemented by the InfoNCE loss as follows: